OpenStudy (anonymous):

How do you find all the solutions to sin^2x + sinx = 0 Is it by factoring? since I got (sinx)^2+sinx= 0

OpenStudy (jamesj):

...hence sin x ( sin x + 1) = 0 This means that sin x = 0 or sin x + 1 = 0

OpenStudy (anonymous):

@JamesJ those are the only solutions? (sinx)^2+sinx=0 and sinx(sinx+1)=0